OpenStudy (anonymous):

A 20 N horizontal force pushes on an 8.0 kg mass resting on a horizontal surface. The surface exerts a friction force of 14 N against the motion. The acceleration of the 8.0 kg mass is ____ m/s^2.

OpenStudy (shubhamsrg):

on which planet ?

OpenStudy (anonymous):

you can't do this.

OpenStudy (anonymous):

context ;) is it even falling? the acceleration of a brick on a table is 0. it's not going anywhere

OpenStudy (anonymous):

*assuming earth and the usual reference frames

OpenStudy (anonymous):

Im sorry i have to put the rest of the question up

OpenStudy (anonymous):

@RONNCC does that help

OpenStudy (anonymous):

yes ;)

OpenStudy (anonymous):

@shubhamsrg i had to add the rest of the question

OpenStudy (anonymous):

ehh static friction right?

OpenStudy (anonymous):

sorry. post this on @Physics.

OpenStudy (anonymous):

just how categories work

OpenStudy (anonymous):

yea

OpenStudy (shubhamsrg):

nevermind physics section, this is simple mathematics vector question ;)

OpenStudy (shubhamsrg):

So, 20 N pushes it, 14N opposes the motion, can you tell me the net force acting on the block ?

OpenStudy (anonymous):

Note: Friction Alwayzz Opposes the Relative Motion

OpenStudy (anonymous):

no am i suppose to divide something

OpenStudy (anonymous):

So it is Alwayzz Taken as - ve

OpenStudy (anonymous):

@Yahoo! so what should i do for the first step

OpenStudy (anonymous):

20N is Applied Force and Friction is 14N F total = 20 - 14 = 6N

OpenStudy (anonymous):

and F = ma a = F/m

OpenStudy (anonymous):

so is 6N the answer

OpenStudy (anonymous):

a = 6/8

OpenStudy (anonymous):

@Yahoo! or is 6/8 the answer